# VoiceShop: A Unified Speech-to-Speech Framework for Identity-Preserving Zero-Shot Voice Editing

### 10 Apr 2024 v1

@@ 4.4 Zero-Shot Combined Multi-Attribute Editing
Subjective CMOS Evaluation. We design a modified CMOS study to measure two key qualities of VoiceShop’s multi-attribute editing capabilities: overall editing capabilities, i.e., participants agree that attributes we claim to edit are in fact edited according to their best judgment, and attribute disentanglement, i.e., the modification of any combination of attributes does not alter participants’ perception of other attributes we claim are kept constant.
Table 10: Results of multi-attribute editing subjective evaluation. The accuracy indicates the proportion of participants who correctly completed each task. We denote the number of responses for each task in parentheses.
Task Accuracy (normal-$\uparrow$)
@@
Correctly observe that any attribute has changed, regardless of whether or not other attributes are edited
@@
Objective Evaluation. We first experiment by editing age and gender using out-of-domain speech samples from VCTK by randomly selecting three speech samples for each of the 109 speakers in this dataset, which are also used by NANSY++ (Choi et al., 2022). We use the attribute predictor in §3.3.1 to analyze the age and gender distributions of VCTK samples before and after our attribute editing, with results shown in Figure 6. We demonstrate that our framework achieves competitive age and gender editing strength, despite that our CNF is only trained on weakly-supervised public attribute datasets. For this comparison, we use our VS-EN-L18 backbone model.
@@
Furthermore, rather than performing serial editing of individual attributes, we showcase our framework’s performance on combined multi-attribute editing, whereby arbitrary unseen users can modify their accent, age, and gender simultaneously in a single forward synthesis pass while preserving their timbre. This capability is enabled by our plug-in modular design that allows concurrent editing on both the speaker embedding and content features of our VS-EN-CN-L18 backbone model.
@@ 5 Conclusion
In this work, we presented VoiceShop, a novel speech-to-speech framework that enables the modification of multiple speech attributes while preserving the input speaker’s timbre in both monolingual and cross-lingual settings. It overcomes the limitations of previous models by enabling simultaneous editing of attributes, providing zero-shot capability for out-of-domain speakers, and avoiding timbre leakage that alters the speaker’s perceived identity. It additionally provides a new assortment of methods to tackle the under-explored voice editing task.
@@ 6 Ethical Considerations
As with all generative artificial intelligence systems, the real-world impact and potential for unintended misuse of models like VoiceShop must be considered. While there are beneficial use cases of our framework, such as providing entertainment value or lowering cross-cultural communication barriers by allowing users to speak other languages or accents in their own voice, its zero-shot capabilities could enable a user to generate misleading content with relative ease, such as synthesizing speech in the voice of an individual without their knowledge, presenting a risk of misinformation. This concern is not unique to VoiceShop and motivates continued efforts towards developing robust audio deepfake detection models deployed in tandem with generative speech models (Guo et al., 2024; Zang et al., 2024; Zhang et al., 2023b) or the incorporation of imperceptible watermarks in synthesized audio signals to classify genuine or generated speech (Cao et al., 2023; Chen et al., 2024; Juvela & Wang, 2024; Liu et al., 2023a; Roman et al., 2024), although this subject is outside the scope of this work. In an effort to balance the need for transparent, reproducible, and socially responsible research practices, and due to the proprietary nature of portions of data used in this work, we share the details of our findings here, but do not plan to publicly release the model checkpoints or implementation at this time. The authors do not condone the use of this technology for illegal or malicious purposes.
@@ A.1 Implementation details of conformer-based ASR models
Figure A1: The architecture of our conditional diffusion backbone model, consisting of an ECAPA-TDNN-based speaker encoder (Desplanques et al., 2020) and one-dimensional U-Net (Ronneberger et al., 2015) inspired by Moûsai (Schneider et al., 2023).
We use the open-source ESPnet [^12] https://github.com/espnet/espnet (Watanabe et al., 2018) library to train our ASR model, following the the ESPnet2 recipe with a conformer-based encoder and a transformer-based decoder. We add an additional head to the output of the conformer encoder to calculate logits for a Connectionist Temporal Classification (CTC) (Graves et al., 2006) loss. The encoder and decoder are connected with cross attention, as in the standard transformer network design (Vaswani et al., 2017). We optimize the model by jointly minimizing CTC and cross entropy (CE) loss terms on the output of the transformer decoder, using a hyperparameter $\lambda$ to interpolate the weights of the two terms. Given paired ASR training data of speech and corresponding text $(X,Y)$, we define the training objective as follows:
@@
Since their debut, conformers (Gulati et al., 2020) have been widely adopted in the ASR community. Compared to conventional transformer encoders, conformers add a convolutional module between the multi-head self-attention and feed-forward module in each computational block, providing the ability to model the local interactions between speech representations besides the global modeling ability enabled by multi-head self-attention. Thus, to learn better semantic representations for VoiceShop, we adopt a conformer-based encoder for our ASR model.
We train monolingual and bilingual versions of the ASR model, referred to as ASR-EN and ASR-EN-CN respectively, such that the former only transcribes English speech and the latter transcribes both English and Mandarin speech. The conformer encoder of both configurations consist of 16 blocks and the transformer decoder respectively contains 6 and 4 blocks in ASR-EN and ASR-EN-CN. ASR-EN uses a hidden dimension of 768 for both encoder and decoder modules, containing approximately 220M learnable parameters, whereas ASR-EN-CN sets this value to 512 with roughly 140M total parameters. Both models set $\lambda$ in Equation 11 to 0.1 and apply SpecAugment (Park et al., 2019) to improve their generalization ability.
ASR-EN is trained on an in-house dataset consisting of 40K hours of English pairs of audio and corresponding transcriptions provided by human annotators, amounting to 4M utterances, while ASR-EN-CN is trained on an in-house dataset consisting of 60K hours of in-house Mandarin data and 5K hours of in-house English data, totaling 5M utterances. We use a joint vocabulary of 11K unique tokens consisting of English subwords prepared using the SentencePiece tokenizer which implements Byte-Pair Encoding (BPE) and Mandarin characters, and omit Mandarin tokens for the monolingual model. Both models are trained using the Adam optimizer (Kingma & Ba, 2014) on 32 V100 GPUs with a learning rate $5\times 10^{-4}$, 20K warm up steps, and a batch size equivalent to 2 hours of audio data for 40 epochs. ASR-EN achieves a word error rate (WER) of 8.86% on an internal English test set and ASR-EN-CN achieves a character error rate (CER) of 7.4% on an internal Chinese test set in the video domain.
@@ A.2 Implementation details of conditional diffusion backbone model
We modify the text-to-music framework Moûsai (Schneider et al., 2023) for the design of our conditional diffusion backbone model, depicted in Figure A1, and describe its implementation details in this section. We base our code on their open-source implementation [^13] https://github.com/archinetai/audio-diffusion-pytorch to enable better reproducibility.
Given a mel-spectrogram of speech, the diffusion model accepts two conditioning signals: an utterance-level global speaker embedding produced by speaker encoder and a time-varying content embedding produced by a pre-trained ASR or APR model. The speaker encoder is trained jointly with the diffusion model from scratch and adopts the same configuration as the ECAPA-TDNN speaker verification model (Desplanques et al., 2020), whereas the weights of the ASR or APR model remain frozen during training.
@@
We adopt the one-dimensional U-Net (Ronneberger et al., 2015) of Moûsai (Schneider et al., 2023) as the architecture of our diffusion model and follow their terminology of “U-Net items” to describe the various components in each U-Net block which incorporate conditioning signals into the model. Please refer to Figure A2 for a visual depiction of our modified U-Net block design. Each U-Net block sequentially processes data in a series of operations consisting of a one-dimensional convolutional “ResNet item” (He et al., 2015), a “modulation item” where the time step and global speaker embeddings are provided to the model, a “MergeAdd item” where the time-varying content features are supplied, and a “self-attention item.” Note that we remove the “cross-attention item” included in their work, as our framework does not use text, and replace their “inject item” with a simpler piecewise addition operation. The details of each item are depicted in Figure A3.
@@
A noise-injected sample is first processed by the “ResNet item,” after which the global condition is provided to the “modulation item,” where it is used to predict the parameters of an affine transformation applied to the output of the “ResNet item.” The local condition is then passed to the “MergeAdd item,” where it is added directly to the output of the “modulation item.” Finally, the sample is passed through a linear projection layer, which predicts query, key, and value tensors received by a standard multi-head self-attention layer (Vaswani et al., 2017). The residual connections of the U-Net are applied using a “MergeConcat” operation depicted in Figure A5.
@@ A.3 Implementation details of vocoder
We adopt HiFi-GAN (Kong et al., 2020a) as our vocoder for its low latency and competitive performance against autoregressive alternatives such as WaveNet (van den Oord et al., 2016), but with key modifications to its architecture and objective function. To address occasional, undesirable glitches in speech synthesized by pre-trained instances of HiFi-GAN, we add a learnable sinusoidal activation module inspired by the source-filter theory of speech production (Fant, 1971), which receives mel-spectrograms of speech and predicts its corresponding excitation signal representing the fundamental frequency ($F_{0}$) contour of an utterance to be provided as extra conditioning signals to the convolutional blocks of the generator and discriminator modules. We make use of the originally proposed multi-period and multi-scale discriminators, adding integrated sub-band coding filters to reduce aliasing, and reduce the receptive field of HiFi-GAN’s convolutional layers to correspond to 7 frames for spectrograms with a hop size of 240.
We employ HiFi-GAN’s mel-spectrogram regression and feature matching loss terms, defined as follows:
@@
such that $\phi$ is the function that converts time-domain waveforms to their mel-spectrogram representations. Furthermore, we add the energy, time, and phase loss terms proposed by TFGAN (Tian et al., 2020), which we find assists the model converge relatively earlier than when omitted, defined as:
@@
such that $x$ and $\hat{x}$ correspond respectively to the ground truth and synthesized audio signals, $\Delta$ refers to the first-difference estimator, and $\|\cdot\|_{1}$ is the $L^{1}$ Frobenius norm. We train the sinsusoidal excitation module using MAE on the ground truth and predicted $F_{0}$ contours, $s$ and $\hat{s}$, respectively:
@@
such that $\lambda_{1}=1,\lambda_{2}=1,\lambda_{3}=100,\lambda_{4}=200,\lambda_{5}=100$, and $\lambda_{6}=1$. We train the vocoder using the Adam optimizer (Kingma & Ba, 2014) for the generator and all discriminators, each with a learning rate of $1\times 10^{-3}$, providing the generator with a warm up of 2K iterations. The model is trained for 2M iterations on 8 V100 GPUs using a batch size of 12 samples on roughly 409 hours of in-house multilingual speech data consisting primarily of English and Mandarin speech, and to a lesser extent additional languages, such as Japanese and Indonesian, after which it robustly generates 24kHz audio.
@@ A.4 Augmenting non-parallel speech into parallel “timbre-matched” datasets
Our proposed bottleneck-to-bottleneck (BN2BN) modeling approach necessitates the use of a parallel training dataset consisting of audio samples of individuals speaking the same content in a variety of accents or speech styles. The collection of recorded speech that meets such requirements is typically understood to be preventatively difficult, time-consuming, and costly in existing literature. One apparent drawback of manually recording such a dataset, apart from other practical concerns, is that it requires a single person to fluently speak in a variety of accents, most of which will not be native to the speaker, which can result in inaccurate depictions of the accents, when instead it would be preferable to collect the speech of native speakers of the various accents to achieve faithful accent and speech style conversion. By leveraging the rapid progress of neural TTS and VC modeling available in a wide array of accents, we demonstrate that acquiring such a dataset can be straightforward, efficient, and cost-effective. In fact, while it is not necessary, our proposed approach can be trained entirely on synthetically generated speech.
@@
A.5 Differences to “BN2BN” module proposed by Zhang et al. (2022a)
In a separate and independent work, Zhang et al. (2022a) also propose the use of a learnable module referred to as “BN2BN” to achieve accent transfer within a TTS framework named AccentSpeech. While both works share a similar intuition of training a BN2BN model to map between the bottleneck features of utterances in a source accent and those of utterances in a target accent, we distinguish our approach from theirs by noting the following key differences in our proposed methods.
@@
Since the decoder is trained on the high-fidelity recordings of a single speaker, AccentSpeech is capable of synthesizing accented speech in one target timbre, whereas VoiceShop aims to preserve any arbitrary source speaker’s timbre in the conversion process. AccentSpeech uses feed-forward transformer blocks stacked between one-dimensional convolutional layers as the model design of its BN2BN module, while we adopt an attention-based sequence-to-sequence architecture. One disadvantage of this design choice on the part of Zhang et al. (2022a) is that AccentSpeech needs frame-wise time-aligned parallel data between two speakers of different accents, requiring third-party forced alignment tools. Another consequence of this restriction is that duration must be considered irrelevant to accent information. VoiceShop does not require frame-level alignments for parallel speech pairs and allows such pairs to be of different durations. Furthermore, VoiceShop is capable of modeling and preserving the varying rhythmic patterns of native speech across accents and speech styles, which, in our view, is an important factor to the perceived fluency of our model’s output speech.

### 11 Apr 2024 v2

@@ 4.4 Zero-Shot Combined Multi-Attribute Editing
Rather than performing serial editing of individual attributes, we showcase our framework’s performance on combined multi-attribute editing, whereby arbitrary unseen users can modify their accent, age, and gender simultaneously in a single forward synthesis pass while preserving their timbre. This capability is enabled by our plug-in modular design that allows concurrent editing on both the speaker embedding and content features of our VS-EN-CN-L18 backbone model.
Subjective CMOS Evaluation.
We design a modified CMOS study to measure two key qualities of VoiceShop’s multi-attribute editing capabilities: overall editing capabilities, i.e., participants agree that attributes we claim to edit are in fact edited according to their best judgment, and attribute disentanglement, i.e., the modification of any combination of attributes does not alter participants’ perception of other attributes we claim are kept constant.
Table 9: Results of multi-attribute editing subjective CMOS evaluation. The accuracy indicates the proportion of participants who correctly completed each task. We denote the number of responses for each task in parentheses. Some tasks have overlapping samples.
Category Task Accuracy (normal-$\uparrow$)
Correctly observe when an attribute has been edited
@@
Correctly observe that any attribute has changed, regardless of which attributes are edited
@@
Correctly observe when an attribute has not been edited
@@
Objective Evaluation.
We first experiment by editing age and gender using out-of-domain speech samples from VCTK by randomly selecting three speech samples for each of the 109 speakers in this dataset, which are also used by NANSY++ (Choi et al., 2022). We use the attribute predictor in §3.3.1 to analyze the age and gender distributions of VCTK samples before and after our attribute editing, with results shown in Figure 6. We demonstrate that our framework achieves competitive age and gender editing strength, despite that our CNF is only trained on weakly-supervised public attribute datasets. For this comparison, we use our VS-EN-L18 backbone model.
@@ 5 Conclusion
In this work, we presented VoiceShop, a novel speech-to-speech framework that enables the modification of multiple speech attributes while preserving the input speaker’s timbre in both monolingual and cross-lingual settings. It overcomes the limitations of previous models by enabling simultaneous editing of attributes, providing zero-shot capability for out-of-domain speakers, and avoiding timbre leakage that alters the speaker’s unedited attributes in voice editing tasks. It additionally provides a new assortment of methods to tackle the under-explored voice editing task.
@@ 6 Ethical Considerations
As with all generative artificial intelligence systems, the real-world impact and potential for unintended misuse of models like VoiceShop must be considered. While there are beneficial use cases of our framework, such as providing entertainment value or lowering cross-cultural communication barriers by allowing users to speak other languages or accents in their own voice, its zero-shot capabilities could enable a user to generate misleading content with relative ease, such as synthesizing speech in the voice of an individual without their knowledge, presenting a risk of misinformation. This concern is not unique to VoiceShop and motivates continued efforts towards developing robust audio deepfake detection models deployed in tandem with generative speech models (Guo et al., 2024; Zang et al., 2024; Zhang et al., 2023b) or the incorporation of imperceptible watermarks in synthesized audio signals to classify genuine or generated speech (Cao et al., 2023; Chen et al., 2024; Juvela and Wang, 2024; Liu et al., 2023a; Roman et al., 2024), although this subject is outside the scope of this work. In an effort to balance the need for transparent, reproducible, and socially responsible research practices, and due to the proprietary nature of portions of data used in this work, we share the details of our findings here, but do not plan to publicly release the model checkpoints or implementation at this time. The authors do not condone the use of this technology for illegal or malicious purposes.
@@ A.1 Implementation details of conformer-based ASR models
Figure A1: The architecture of our conditional diffusion backbone model, consisting of an ECAPA-TDNN-based speaker encoder [Desplanques et al., 2020] and one-dimensional U-Net [Ronneberger et al., 2015] inspired by Moûsai [Schneider et al., 2023].
We use the open-source ESPnet [^12] https://github.com/espnet/espnet [Watanabe et al., 2018] library to train our ASR model, following the ESPnet2 recipe with a conformer-based encoder and a transformer-based decoder. We add an additional head to the output of the conformer encoder to calculate logits for a Connectionist Temporal Classification (CTC) [Graves et al., 2006] loss. The encoder and decoder are connected with cross attention, as in the standard transformer network design [Vaswani et al., 2017]. We optimize the model by jointly minimizing CTC and cross-entropy (CE) loss terms on the output of the transformer decoder, using a hyperparameter $\lambda$ to interpolate the weights of the two terms. Given paired ASR training data of speech and corresponding text $(X,Y)$, we define the training objective as follows:
@@
Since their debut, conformers [Gulati et al., 2020] have been widely adopted in the ASR community. Compared to conventional transformer encoders, conformers add a convolutional module between the multi-head self-attention and feed-forward module in each computational block, providing the ability to model the local interactions between speech representations besides the global modeling ability enabled by multi-head self-attention. Thus, to learn better semantic representations for VoiceShop, we adopt a conformer-based encoder for our ASR model.
We train monolingual and bilingual versions of the ASR model, referred to as ASR-EN and ASR-EN-CN respectively, such that the former only transcribes English speech and the latter transcribes both English and Mandarin speech. The conformer encoder of both configurations consists of 16 blocks and the transformer decoder respectively contains 6 and 4 blocks in ASR-EN and ASR-EN-CN. ASR-EN uses a hidden dimension of 768 for both encoder and decoder modules, containing approximately 220M learnable parameters, whereas ASR-EN-CN sets this value to 512 with roughly 140M total parameters. Both models set $\lambda$ in Equation 12 to 0.1 and apply SpecAugment [Park et al., 2019] to improve their generalization ability.
ASR-EN is trained on an in-house dataset consisting of 40K hours of English pairs of audio and corresponding transcriptions provided by human annotators, amounting to 4M utterances, while ASR-EN-CN is trained on an in-house dataset consisting of 60K hours of in-house Mandarin data and 5K hours of in-house English data, totaling 5M utterances. We use a joint vocabulary of 11K unique tokens consisting of English subwords prepared using the SentencePiece tokenizer which implements Byte-Pair Encoding (BPE) and Mandarin characters, and omits Mandarin tokens for the monolingual model. Both models are trained using the Adam optimizer [Kingma and Ba, 2014] on 32 V100 GPUs with a learning rate $5\times 10^{-4}$, 20K warm up steps, and a batch size equivalent to 2 hours of audio data for 40 epochs. ASR-EN achieves a word error rate (WER) of 8.86% on an internal English test set and ASR-EN-CN achieves a character error rate (CER) of 7.4% on an internal Chinese test set in the video domain.
@@ A.2 Implementation details of conditional diffusion backbone model
We modify the text-to-music framework Moûsai [Schneider et al., 2023] for the design of our conditional diffusion backbone model, depicted in Figure A1, and describe its implementation details in this section. We base our code on their open-source implementation [^13] https://github.com/archinetai/audio-diffusion-pytorch to enable better reproducibility.
Given a mel-spectrogram of speech, the diffusion model accepts two conditioning signals: an utterance-level global speaker embedding produced by a speaker encoder and a time-varying content embedding produced by a pre-trained ASR or APR model. The speaker encoder is trained jointly with the diffusion model from scratch and adopts the same configuration as the ECAPA-TDNN speaker verification model [Desplanques et al., 2020], whereas the weights of the ASR or APR model remain frozen during training.
@@
We adopt the one-dimensional U-Net [Ronneberger et al., 2015] of Schneider et al. [2023] as the architecture of our diffusion model and follow their terminology of “U-Net items” to describe the various components in each U-Net block which incorporate conditioning signals into the model. Please refer to Figure A2 for a visual depiction of our modified U-Net block design. Each U-Net block sequentially processes data in a series of operations consisting of a one-dimensional convolutional “ResNet item” [He et al., 2015], a “modulation item” where the time step and global speaker embeddings are provided to the model, a “MergeAdd item” where the time-varying content features are supplied, and a “self-attention item.” Note that we remove the “cross-attention item” included in their work, as our framework does not use text, and replace their “inject item” with a simpler piecewise addition operation. The details of each item are depicted in Figure A3.
@@
A noise-injected sample is first processed by the “ResNet item,” after which the global condition is provided to the “modulation item,” where it is used to predict the parameters of an affine transformation applied to the output of the “ResNet item.” The local condition is then passed to the “MergeAdd item,” where it is added directly to the output of the “modulation item.” Finally, the sample is passed through a linear projection layer, which predicts query, key, and value tensors received by a standard multi-head self-attention layer [Vaswani et al., 2017]. The residual connections of the U-Net are applied using a “MergeConcat” operation depicted in Figure A5.
@@ A.3 Implementation details of vocoder
We adopt HiFi-GAN [Kong et al., 2020a] as our vocoder for its low latency and competitive performance against autoregressive alternatives such as WaveNet [van den Oord et al., 2016], but with key modifications to its architecture and objective function. To address occasional, undesirable glitches in speech synthesized by pre-trained instances of HiFi-GAN, we add a learnable sinusoidal activation module inspired by the source-filter theory of speech production [Fant, 1971], which receives mel-spectrograms of speech and predicts its corresponding excitation signal representing the fundamental frequency ($F_{0}$) contour of an utterance to be provided as extra conditioning signals to the convolutional blocks of the generator and discriminator modules. We make use of the originally proposed multi-period and multi-scale discriminators, adding integrated sub-band coding filters to reduce aliasing, and reduce the receptive field of HiFi-GAN’s convolutional layers to correspond to 7 frames for spectrograms with a hop size of 240.
We employ HiFi-GAN’s mel-spectrogram regression and feature-matching loss terms, as follows:
@@
such that $\phi$ is the function that converts time-domain waveforms to their mel-spectrogram representations. Furthermore, we add the energy, time, and phase loss terms proposed by TFGAN [Tian et al., 2020], which we find assists the model converge relatively earlier than when omitted, defined as:
@@
such that $x$ and $\hat{x}$ correspond respectively to the ground truth and synthesized audio signals, $\Delta$ refers to the first-difference estimator, and $\|\cdot\|_{1}$ is the $L^{1}$ Frobenius norm. We train the sinusoidal excitation module using MAE on the ground truth and predicted $F_{0}$ contours, $s$ and $\hat{s}$, respectively:
@@
such that $\lambda_{1}=1,\lambda_{2}=1,\lambda_{3}=100,\lambda_{4}=200,\lambda_{5}=100$, and $\lambda_{6}=1$. We train the vocoder using the Adam optimizer [Kingma and Ba, 2014] for the generator and all discriminators, each with a learning rate of $1\times 10^{-3}$, providing the generator with a warm up of 2K iterations. The model is trained for 2M iterations on 8 V100 GPUs using a batch size of 12 samples on roughly 409 hours of in-house multilingual speech data consisting primarily of English and Mandarin speech, and to a lesser extent additional languages, such as Japanese and Indonesian, after which it robustly generates 24kHz audio.
@@ A.4 Augmenting non-parallel speech into parallel “timbre-matched” datasets
Our proposed bottleneck-to-bottleneck (BN2BN) modeling approach necessitates the use of a parallel training dataset consisting of audio samples of individuals speaking the same content in a variety of accents or speech styles. The collection of recorded speech that meets such requirements is typically understood to be preventatively difficult, time-consuming, and costly in existing literature. One apparent drawback of manually recording such a dataset, apart from other practical concerns, is that it requires a single person to fluently speak in a variety of accents, most of which will not be native to the speaker, which can result in inaccurate depictions of the accents when instead it would be preferable to collect the speech of native speakers of the various accents to achieve faithful accent and speech style conversion. By leveraging the rapid progress of neural TTS and VC modeling available in a wide array of accents, we demonstrate that acquiring such a dataset can be straightforward, efficient, and cost-effective. In fact, while it is not necessary, our proposed approach can be trained entirely on synthetically generated speech.
@@
A.5 Differences to “BN2BN” module proposed by Zhang et al. [2022a]
In a separate and independent work, Zhang et al. [2022a] also proposes the use of a learnable module referred to as “BN2BN” to achieve accent transfer within a TTS framework named AccentSpeech. While both works share a similar intuition of training a BN2BN model to map between the bottleneck features of utterances in a source accent and those of utterances in a target accent, we distinguish our approach from theirs by noting the following key differences in our proposed methods.
@@
Since the decoder is trained on the high-fidelity recordings of a single speaker, AccentSpeech is capable of synthesizing accented speech in one target timbre, whereas VoiceShop aims to preserve any arbitrary source speaker’s timbre in the conversion process. AccentSpeech uses feed-forward transformer blocks stacked between one-dimensional convolutional layers as the model design of its BN2BN module, while we adopt an attention-based sequence-to-sequence architecture. One disadvantage of this design choice on the part of Zhang et al. [2022a] is that AccentSpeech needs frame-wise time-aligned parallel data between two speakers of different accents, requiring third-party forced alignment tools. Another consequence of this restriction is that duration must be considered irrelevant to accent information. VoiceShop does not require frame-level alignments for parallel speech pairs and allows such pairs to be of different durations. Furthermore, VoiceShop is capable of modeling the varying rhythmic patterns of native speech across accents and speech styles, which, in our view, is an important factor to the perceived fluency of our model’s output speech.